OpenStudy (a1234):

Solve algebraically over the interval 0 <= x <= 2pi. Round to the nearest hundredth. csc^2 x = 5 x = 0.46 and -0.46

OpenStudy (danjs):

csc is 1/sin 1/(sin x)^2 = 5

OpenStudy (danjs):

recall \[\sqrt{u^2}=\left| u \right|= +u ~~if ~~u \ge 0~~~~~or~~~~~-u~~if ~~u<0\]

OpenStudy (a1234):

So how you solve this?